### v1

@@ 5 Recommendations for Practice
Our review of the empirical literature employing BD designs showed that the overwhelming majority of studies focus exclusively on the identification, estimation and inference for the BATE using pooling-based methods (Section 3). This scalar parameter captures a density-weighted average of potentially heterogeneous treatment effects along the assignment boundary. While modern empirical work has recognized the importance of localization and flexible local parametrization, the current literature often exhibits three key limitations: (i) local regression specifications frequently omit the interaction between the treatment assignment variable $T_{i}$ and the flexible regression terms $\mathbf{r}_{p}(D_{i})$ or $\mathbf{r}_{p}(\mathbf{X}_{i})$, risking significant misspecification bias; (ii) the bandwidth $h$ is often chosen in an ad hoc manner; and (iii) the role of misspecification bias from a “large” bandwidth, such as an MSE-optimal choice, is often ignored, which can lead to over-rejection of the null hypothesis. Cattaneo-Titiunik-Yu_2025_BDD-Pooling provide a formal study of pooling-based methods that explicitly addresses these limitations. Moving forward, we recommend that empirical researchers implementing pooling-based methods always use specifications (6) or (7), select an MSE-optimal bandwidth, and conducting inference using robust bias-corrected methods.
More broadly, as noted by Cattaneo-Titiunik-Yu_2025_BDD-Distance; Cattaneo-Titiunik-Yu_2025_BDD-Location, a more general approach to analyzing and interpreting BD designs is to first focus on the BATEC parameter, and then consider other aggregate causal treatment effects as transformations thereof. As summarized in Section 4, distance-based and location-based methods capture the full richness of the BD design by estimating average treatment effects at each point along the assignment boundary. These methods also recover the BATE parameter as a special case, in addition to other interesting causal parameters such as the LBATE. These methods can be implemented using the general-purpose R package rd2d available at https://rdpackages.github.io/, and discussed in Cattaneo-Titiunik-Yu_2025_rd2d. In particular, location-based methods are the most robust and general approach; we therefore recommend them for future empirical work leveraging BD designs.
@@
holmes1998effect Manufacturing activity Right-to-work policies Lat, Lon Euclidean (1), (2), (6) Manual Yes
black1999better House prices School quality Lat, Lon Euclidean (1), (2) Manual No
kane2003school House prices School quality Lat, Lon Euclidean (1), (2) Manual No
kane2006school Housing prices School quality Lat, Lon Euclidean (1), (2) Manual No
bayer2007unified House prices School quality Lat, Lon Euclidean (1), (2) Manual No
lalive2008extended Unemployment duration Unemployment benefits Lat, Lon Driving time (6) Manual No
dell2010persistent Consumption, childhood stunting Forced mining labor Lat, Lon Euclidean (3) MSE No
ou2010leave High school dropout Failure of high school exit exam Mathematics score, Language score Euclidean (8) Manual No
dube2010minimum Employment Increase in minimum wage Lat, Lon County adjacency (1), (2) Manual No
grout2011land Property values Land-use regulations Lat, Lon Euclidean (8) Manual Yes
robinson2011evaluating English proficiency English proficiency reclassification Scores on five language tests Euclidean Binding score MSE No
eugster2011demand Demand for social insurance Culture Lat, Lon Driving kilometers (6) Manual No
hinnerich2014democracy Political regimes Income redistribution Population in two separate years Euclidean (8) MSE No
ferwerda2014political Resistance activity Devolution of governing authority Lat, Lon Euclidean (6) MSE, Manual No
snider2015barriers Airline industry outcomes Access to airport facilities granted to new airlines Share of passengers in top two carriers in origin and destination city (1), (11) MSE, Manual Yes
barone2015telecracy Vote shares Switch from analog to digital TV Lat, Lon Euclidean (5) Manual No
michalopoulos2016long Civil conflict Ethnic partitioning Lat, Lon Euclidean (1), (2) Manual No
egger2015impact Foreign investment Controlled foreign company rules Tax rate, Income measures NA (7) $L=1$ MSE No
Keele-Titiunik_2015_PA Voter turnout Political TV advertisement Lat, Lon Chordal (10) MSE Yes
macdonald2016effect Crime Policing Lat, Lon Euclidean (6) MSE No
evans2017smart STEM major Financial incentive GPA, Family contribution score Euclidean (8) MSE No
eugster2017culture Unemployment Culture Lat, Lon Road distance (6) Manual No
kumar2018restrictions Mortgage default Limits on home equity borrowing Lat, Lon Euclidean (4), (3) Manual No
@@
dell2018nation Public goods, political attitudes War strategies Lat, Lon Euclidean (3) Manual No
@@
clinton2018politics Voter registration and turnout Medicaid expansion Lat, Lon Euclidean (6) Manual No
@@
spenkuch2018political Voter turnout Television advertisement Lat, Lon Euclidean (2), (6) MSE, Manual No
@@
de2018agents Party vote shares Political alignment Lat, Lon Euclidean (3) Manual No
eugster2019culture Tax competition Culture Lat, Lon Road (1) in DID Manual No
@@
giuntella2019sunset Sleep duration, Health outcomes Timing of natural light Lat, Lon Euclidean (6) MSE, Manual No
@@
johnson2019effects High School graduation English learner classification Scores on two language tests Euclidean Binding score MSE No
@@
velez2019tuning Voter turnout Spanish-language television station Lat, Lon Euclidean (1), (2), (10) MSE, Manual Yes
dupraz2019french Schooling Colonial legacies Lat, Lon Euclidean (6), (3) MSE No
moscona2020segmentary Conflict Segmentary lineage organization Lat, Lon Euclidean (6) Manual No
ambrus2020loss Property value Cholera epidemic Lat, Lon Walking (6) $L=1$ MSE, Manual No
@@
dell2020development Schools, education levels Sugar cultivation for Dutch cultivation system Lat, Lon Euclidean (3) Manual No
he2020watering Pollution Water quality monitoring Lat, Lon Euclidean (6) $L=1$ MSE No or NA
ito2020willingness Willingness to pay for clean air Lat, Lon Euclidean, road (6) MSE, Manual No
wuepper2020countriesErosion Soil erosion Country borders Lat, Lon Euclidean (6) MSE No
albertus2020land Conflict Land reform Lat, Lon Euclidean (6) MSE, Manual No
@@
wuepper2020countriesPollution Crop yield gaps, nitrogen pollution Country borders Lat, Lon Euclidean (6) MSE No
@@
schafer2020time Voter turnout Time zones Lat, Lon Chordal (6) MSE, Manual No
@@
letsa2020mechanisms Attitudes towards local power Colonial legacies Lat, Lon Euclidean (3) Manual No
@@
LondonoVelezRodriguezSanchez_2020_AEJ College enrollment Financial subsidy High school exit exam score, wealth index Euclidean (8) MSE No
gonzalez2021cell Election fraud Cell phone coverage Lat, Lon Euclidean (6), (7) MSE Yes
ayres2021environmental Land value Adjudication of groundwater rights Lat, Lon Euclidean (6) $L=1$ MSE No
laliberte2021long School quality Educational achievement Lat, Lon Euclidean (6) MSE No
@@
aaronson2021effects Characteristics of urban neighborhoods Red lining maps Lat, Lon Euclidean (2) in DID Manual No
@@
appau2021long Agricultural productivity Bombing intensity Lat, Lon Absolute difference between centroid and 17th parallel north latitude (4) (first stage) Manual No
@@
lowes2021concessions Education, health, wealth Concession to extract rubber Lat, Lon Euclidean (6) MSE, Manual No
sides2022effect Election outcomes Television advertisement Lat, Lon Euclidean (2) Manual No
@@
zheng2022valuation House prices Charter school entry Lat, Lon Euclidean Event study in sample near boundary Manual No
@@
dehdari2022origins Regional identity Actions of nation states Lat, Lon Euclidean (6) MSE, Manual No
@@
mendez2022multinationals Living standards Land concession to multinational firm Lat, Lon Euclidean (3) Manual No
jones2022factor Profits, irrigation efficiency Irrigation Lat, Lon Euclidean (6) Manual No
@@
henn2023complements State capacity Remoteness from administrative headquarters Lat, Lon Euclidean (6) Manual No
@@
jones2022effects Enrollment, graduation Tuition scholarship High school GPA, SAT/ACT score NA (7) Manual No
@@
dai2022effects Lifestyle outcomes Hypertension diagnosis Systolic and Diastolic blood pressure Euclidean (8) MSE No
@@
salti2022impact Economic outcomes Cash assistance Two scores in means testing formula NA (8) MSE No
@@
castro2022effect Teacher retention, student learning Recruitment bonus Population, distance to capital Euclidean (8) MSE No
@@
moussa2022impact Children health outcomes Cash assistance Two scores in means testing formula NA (8) MSE No
@@
mangonnet2022playing Protected area designation Political alignment Lat, Lon Euclidean (2) Manual No
baragwanath2023collective Forest growth Property rights Lat, Lon Euclidean (6) MSE No
@@
larsen2024long Educational attainment Grade retention Mathematics and English Language Arts test scores Euclidean (8) MSE No
woller2023cost Vote buying Cost of voting Lat, Lon Euclidean (4) Manual No
@@
paulsen2023foundations Voter turnout, economic outcomes State funding for common schools Lat, Lon Euclidean (3) Manual No
@@
prillaman2023strength Political participation Participation in women-only credit groups Lat, Lon Euclidean (2), (3) Manual No
mcalexander2023international Rebel actions UN partition line of Palestine Lat, Lon Euclidean (6) $L=1$ MSE No
@@
murphy2023dual Special education placement English learner designation Various scores of language assessment NA Binding score MSE No
@@
bjerre2024attendance School enrollment School change Lat, Lon Euclidean (2) in DID Manual No
@@
kampfen2024heterogeneous Blood pressure Blood pressure diagnosis Systolic and Diastolic blood pressure NA Binding score MSE No
wuepper2024public Deforestation Country borders Lat, Lon Euclidean (6) MSE No
cox2024bound Voter mobilization Social networks Lat, Lon Fastest driving (1) NA No
@@
grasse2024state Poverty, economic development Mass repression Lat, Lon Euclidean (5) Manual No
jardim2024local Wages, employment Increase in minimum wage Lat, Lon Driving time (1), (2) Manual No
ring2025wealth Savings Wealth tax assessment Lat, Lon Euclidean (6) Manual No
yamagishi2025persistent Land price Buraku neighborhoods Lat, Lon Euclidean (1), (6) $L=1$ MSE, Manual No
loumeau2025regional Commuting flows, residential decisions Departmental borders Lat, Lon Euclidean (5) MSE No
boix2025political Jewish National identity Political emancipation Lat, Lon Euclidean (6) MSE No
muller2025building Share of predominant ethnic group Administrative borders Lat, Lon Euclidean (6) Manual No

### v2

@@ 5 Recommendations for Practice
Our review of the empirical literature employing BD designs showed that the overwhelming majority of studies focuses exclusively on identification, estimation, and inference for the BATE using pooling-based methods (Section 3). This scalar parameter captures a density-weighted average of potentially heterogeneous treatment effects along the assignment boundary. While modern empirical work has recognized the importance of localization and flexible local parametrization, the current literature often exhibits three key limitations: (i) local regression specifications frequently omit the interaction between the treatment assignment variable $T_{i}$ and the flexible regression terms $\mathbf{r}_{p}(D_{i})$ or $\mathbf{r}_{p}(\mathbf{X}_{i})$, risking significant misspecification bias; (ii) the bandwidth $h$ is often chosen in an ad-hoc manner; and (iii) the role of misspecification bias from a large bandwidth, such as an MSE-optimal choice, is often ignored, which can lead to over-rejection of the null hypothesis. Cattaneo et al. (2026) provide a formal study of pooling-based methods that explicitly addresses these limitations. Moving forward, we recommend that empirical researchers implementing pooling-based methods use specifications (6) or (7), select an MSE-optimal bandwidth, and conduct inference using robust bias-corrected methods.
As noted by Cattaneo et al. (2025a, b), a more general approach to analyzing and interpreting BD designs is to first focus on the BATEC parameter, and then consider other aggregate causal treatment effects as transformations thereof. As summarized in Section 4, distance-based and location-based methods capture the full richness of the BD design by estimating average treatment effects at each point along the assignment boundar, and also recover the BATE parameter as a special case. These methods can be implemented using the general-purpose R package rd2d available at https://rdpackages.github.io/, and discussed in Cattaneo et al. (2025c). In particular, location-based methods are the most robust and general approach; we therefore recommend them for future empirical work analyzing BD designs.
Finally, our discussion and literature review have focused on BD designs where the running variable is explicitly bivariate, such as geographic location or test scores in two exams. A more general view of BD designs also includes univariate RD designs with multiple cutoffs. As discussed by Cattaneo et al. (2016a) and Cattaneo et al. (2024), a univariate RD design with multiple cutoffs can be recast as a BD design where one dimension of the bivariate score is the original univariate score and the other is the cutoff variable, and the boundary is the set where these dimensions are equal—for an empirical example where multiple cutoffs induce a known boundary see the seminal paper by Angrist and Lavy (1999). In this sense, with appropriate modifications, our discussion of BD design methodology also applies to RD designs with multiple cutoffs.
@@
Holmes (1998) Manufacturing activity Right-to-work policies Lat, Lon Euclidean (1), (2), (6) Manual Yes
Black (1999) House prices School quality Lat, Lon Euclidean (1), (2) Manual No
Kane et al. (2003) House prices School quality Lat, Lon Euclidean (1), (2) Manual No
Kane et al. (2006) Housing prices School quality Lat, Lon Euclidean (1), (2) Manual No
Bayer et al. (2007) House prices School quality Lat, Lon Euclidean (1), (2) Manual No
Lalive (2008) Unemployment duration Unemployment benefits Lat, Lon Driving time (6) Manual No
Dell (2010) Consumption, childhood stunting Forced mining labor Lat, Lon Euclidean (3) MSE No
Ou (2010) High school dropout Failure of high school exit exam Mathematics score, Language score Euclidean (8) Manual No
Dube et al. (2010) Employment Increase in minimum wage Lat, Lon County adjacency (1), (2) Manual No
Grout et al. (2011) Property values Land-use regulations Lat, Lon Euclidean (8) Manual Yes
Robinson (2011) English proficiency English proficiency reclassification Scores on five language tests Euclidean Binding score MSE No
Eugster et al. (2011) Demand for social insurance Culture Lat, Lon Driving kilometers (6) Manual No
Hinnerich and Pettersson-Lidbom (2014) Political regimes Income redistribution Population in two separate years Euclidean (8) MSE No
Ferwerda and Miller (2014) Resistance activity Devolution of governing authority Lat, Lon Euclidean (6) MSE, Manual No
Snider and Williams (2015) Airline industry outcomes Access to airport facilities granted to new airlines Share of passengers in top two carriers in origin and destination city (1), (11) MSE, Manual Yes
Barone et al. (2015) Vote shares Switch from analog to digital TV Lat, Lon Euclidean (5) Manual No
Michalopoulos and Papaioannou (2016) Civil conflict Ethnic partitioning Lat, Lon Euclidean (1), (2) Manual No
Egger and Wamser (2015) Foreign investment Controlled foreign company rules Tax rate, Income measures NA (7) $L=1$ MSE No
Keele and Titiunik (2015) Voter turnout Political TV advertisement Lat, Lon Chordal (10) MSE Yes
MacDonald et al. (2016) Crime Policing Lat, Lon Euclidean (6) MSE No
Evans (2017) STEM major Financial incentive GPA, Family contribution score Euclidean (8) MSE No
Eugster et al. (2017) Unemployment Culture Lat, Lon Road distance (6) Manual No
Kumar (2018) Mortgage default Limits on home equity borrowing Lat, Lon Euclidean (4), (3) Manual No
@@
Dell and Querubin (2018) Public goods, political attitudes War strategies Lat, Lon Euclidean (3) Manual No
@@
Clinton and Sances (2018) Voter registration and turnout Medicaid expansion Lat, Lon Euclidean (6) Manual No
@@
Spenkuch and Toniatti (2018) Voter turnout Television advertisement Lat, Lon Euclidean (2), (6) MSE, Manual No
@@
De Kadt and Larreguy (2018) Party vote shares Political alignment Lat, Lon Euclidean (3) Manual No
Eugster and Parchet (2019) Tax competition Culture Lat, Lon Road (1) in DID Manual No
@@
Giuntella and Mazzonna (2019) Sleep duration, Health outcomes Timing of natural light Lat, Lon Euclidean (6) MSE, Manual No
@@
Johnson (2019) High School graduation English learner classification Scores on two language tests Euclidean Binding score MSE No
@@
Velez and Newman (2019) Voter turnout Spanish-language television station Lat, Lon Euclidean (1), (2), (10) MSE, Manual Yes
Dupraz (2019) Schooling Colonial legacies Lat, Lon Euclidean (6), (3) MSE No
Moscona et al. (2020) Conflict Segmentary lineage organization Lat, Lon Euclidean (6) Manual No
Ambrus et al. (2020) Property value Cholera epidemic Lat, Lon Walking (6) $L=1$ MSE, Manual No
@@
Dell and Olken (2020) Schools, education levels Sugar cultivation for Dutch cultivation system Lat, Lon Euclidean (3) Manual No
He et al. (2020) Pollution Water quality monitoring Lat, Lon Euclidean (6) $L=1$ MSE No or NA
Ito and Zhang (2020) Willingness to pay for clean air Lat, Lon Euclidean, road (6) MSE, Manual No
Wuepper et al. (2020a) Soil erosion Country borders Lat, Lon Euclidean (6) MSE No
Albertus (2020) Conflict Land reform Lat, Lon Euclidean (6) MSE, Manual No
@@
Wuepper et al. (2020b) Crop yield gaps, nitrogen pollution Country borders Lat, Lon Euclidean (6) MSE No
@@
Schafer and Holbein (2020) Voter turnout Time zones Lat, Lon Chordal (6) MSE, Manual No
@@
Letsa and Wilfahrt (2020) Attitudes towards local power Colonial legacies Lat, Lon Euclidean (3) Manual No
@@
Londoño-Vélez et al. (2020) College enrollment Financial subsidy High school exit exam score, wealth index Euclidean (8) MSE No
Gonzalez (2021) Election fraud Cell phone coverage Lat, Lon Euclidean (6), (7) MSE Yes
Ayres et al. (2021) Land value Adjudication of groundwater rights Lat, Lon Euclidean (6) $L=1$ MSE No
Laliberté (2021) School quality Educational achievement Lat, Lon Euclidean (6) MSE No
@@
Aaronson et al. (2021) Characteristics of urban neighborhoods Red lining maps Lat, Lon Euclidean (2) in DID Manual No
@@
Appau et al. (2021) Agricultural productivity Bombing intensity Lat, Lon Absolute difference between centroid and 17th parallel north latitude (4) (first stage) Manual No
@@
Lowes and Montero (2021) Education, health, wealth Concession to extract rubber Lat, Lon Euclidean (6) MSE, Manual No
Sides et al. (2022) Election outcomes Television advertisement Lat, Lon Euclidean (2) Manual No
@@
Zheng (2022) House prices Charter school entry Lat, Lon Euclidean Event study in sample near boundary Manual No
@@
Dehdari and Gehring (2022) Regional identity Actions of nation states Lat, Lon Euclidean (6) MSE, Manual No
@@
Méndez and Van Patten (2022) Living standards Land concession to multinational firm Lat, Lon Euclidean (3) Manual No
Jones et al. (2022a) Profits, irrigation efficiency Irrigation Lat, Lon Euclidean (6) Manual No
@@
Henn (2023) State capacity Remoteness from administrative headquarters Lat, Lon Euclidean (6) Manual No
@@
Jones et al. (2022b) Enrollment, graduation Tuition scholarship High school GPA, SAT/ACT score NA (7) Manual No
@@
Dai et al. (2022) Lifestyle outcomes Hypertension diagnosis Systolic and Diastolic blood pressure Euclidean (8) MSE No
@@
Salti et al. (2022) Economic outcomes Cash assistance Two scores in means testing formula NA (8) MSE No
@@
Castro and Esposito (2022) Teacher retention, student learning Recruitment bonus Population, distance to capital Euclidean (8) MSE No
@@
Moussa et al. (2022) Children health outcomes Cash assistance Two scores in means testing formula NA (8) MSE No
@@
Mangonnet et al. (2022) Protected area designation Political alignment Lat, Lon Euclidean (2) Manual No
Baragwanath et al. (2023) Forest growth Property rights Lat, Lon Euclidean (6) MSE No
@@
Larsen and Valant (2024) Educational attainment Grade retention Mathematics and English Language Arts test scores Euclidean (8) MSE No
Woller et al. (2023) Vote buying Cost of voting Lat, Lon Euclidean (4) Manual No
@@
Paulsen et al. (2023) Voter turnout, economic outcomes State funding for common schools Lat, Lon Euclidean (3) Manual No
@@
Prillaman (2023) Political participation Participation in women-only credit groups Lat, Lon Euclidean (2), (3) Manual No
McAlexander (2023) Rebel actions UN partition line of Palestine Lat, Lon Euclidean (6) $L=1$ MSE No
@@
Murphy and Johnson (2023) Special education placement English learner designation Various scores of language assessment NA Binding score MSE No
@@
Bjerre-Nielsen and Gandil (2024) School enrollment School change Lat, Lon Euclidean (2) in DID Manual No
@@
Kämpfen and Mosca (2024) Blood pressure Blood pressure diagnosis Systolic and Diastolic blood pressure NA Binding score MSE No
Wuepper et al. (2024) Deforestation Country borders Lat, Lon Euclidean (6) MSE No
Cox et al. (2024) Voter mobilization Social networks Lat, Lon Fastest driving (1) NA No
@@
Grasse (2024) Poverty, economic development Mass repression Lat, Lon Euclidean (5) Manual No
Jardim et al. (2024) Wages, employment Increase in minimum wage Lat, Lon Driving time (1), (2) Manual No
Ring (2025) Savings Wealth tax assessment Lat, Lon Euclidean (6) Manual No
Yamagishi and Sato (2025) Land price Buraku neighborhoods Lat, Lon Euclidean (1), (6) $L=1$ MSE, Manual No
Loumeau (2025) Commuting flows, residential decisions Departmental borders Lat, Lon Euclidean (5) MSE No
Boix (2025) Jewish National identity Political emancipation Lat, Lon Euclidean (6) MSE No
Müller-Crepon (2025) Share of predominant ethnic group Administrative borders Lat, Lon Euclidean (6) Manual No